Hector is performing a chemistry experiment that requires 140 millimeters of a 30% copper sulfate solution. He has a 25% copper sulfate solution and a 60% copper sulfate solution. How many milliliters of each solution should he mix to obtain the needed solution?: Hector is performing a chemistry experiment that requires 140 millimeters of a 30% copper sulfate solution. He has a 25% copper sulfate solution and a 60% copper sulfate solution. How many milliliters of each solution should he mix to obtain the needed solution?
Answer by ankor@dixie-net.com(4683)

Hector is performing a chemistry experiment that requires 140 millimeters of
a 30% copper sulfate solution. He has a 25% copper sulfate solution and a 60%
copper sulfate solution. How many milliliters of each solution should he mix
to obtain the needed solution?
:
Let x = amt of 60% solution
:
It states that total resulting solution should be 140 ml, therefore
(140-x) = amt of 25% solution required
:
The equation:
.60x + .25(140-x) = .30(140)
.60x + 35 - .25x = 42
.60x - .25x = 42 - 35
.35x = 7
x = 7/.35
x = 20 ml of 60% solution
and
140 - 20 = 120 ml of 25% solution
;
:
Check:
.6(20) + .25(120) = .3(140)
12 + 30 = 42